Sunora corp bond is paying 10% interest for 20 years on a semiannual basis. assume interest rates in the market (yield to Maturity) increase from 9 to 12 percent.

a. What is the bond price at 9 percent?

b. What is the bond price at 12 percent?

c. What would be your percentage return on the investment if you bought the bond when rates were 9 percent and sold when rates were 12 percent?

Include financial calculator steps, including the keys pressed on the calculator to solve each step.
Answer #1

a)
FV = 1000
N = 20 * 2 = 40
I/Y = 9% / 2 = 4.5%
PMT = 1000 * 10% / 2 = 50

CPT PV = $1092.01

Bond price = $1092.01


b)
FV = 1000
N = 20 * 2 = 40
I/Y = 12% / 2 = 6%
PMT = 1000 * 10% / 2 = 50

CPT PV = $849.54

Bond price = $849.54


c)
Percentage return = (849.54 - 1092.01) / 1092.01 = -22.20%